Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night report 2017-11-28

Observers: Z.Kostrzewa-Rutkowska, K.Lopez
TAT: J. Chen, M. Fraser, J. Sollerman
Support team: P. Jonker, G. Cannizzaro, F. Onori
Instruments: EFOSC2

Weather conditions: 
evening UT 23:45: clear according to meteo monitor, but some cirrus clouds visible, humidity 42%
morning UT 9:00: thin clouds, humidity 27%

Calibrations:
EFOSC2  bias, arcs daytime, flats evening, morning BVri, arcs morning

LOG:

UT      Target           Type       Filter/Grism ExpTime[s]  Slit   Seeing  Airmass    Comments
00:25   Feige110         STD        gr11/gr13/gr16          1.0/1.5 2.06     1.13
00:50   SMT17scn         Followup   gr13                    1.5     1.97     1.52      Aborted - target not visible at aq image
-------------------
00:58   DDT
-------------------
03:11   ATLAS17mvx       Followup   gr13         2700       1.0     1.31     1.69  
04:07   AT2017imy        Class      gr13          600       1.0     1.21     1.67
04:28   L745a            STD        gr11/gr13/gr16          1.0/1.5 1.26     1.46
04:54   AT2017fzw        Followup   gr11         1800       1.0     0.96     1.07      Image quality fixed
05:13   AT2017fzw        Followup   gr11         1800       1.0     0.95     1.04      Restarted
05:52   AT2017hxz        Followup   gr11         3600       1.0     1.40     1.22
07:00   AT2017ipa        Class      gr13          600       1.0     1.00     1.03
07:20   AT2017fzw        Followup   gr16         1800       1.0     1.02     1.02
08:10   L745a            STD        gr11/gr13/gr16          1.0/1.5 1.20     1.03 

 
Comments